OpenStudy (anonymous):

factoring quadratic form. x^4+x^2-12

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

Let x^2 = y

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how you solve?

OpenStudy (anonymous):

Rewrite your original question with y = x^2

OpenStudy (anonymous):

y^2+y-12?

OpenStudy (anonymous):

Can u do that one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Go on then.

OpenStudy (anonymous):

y(y+1)-12

OpenStudy (anonymous):

(y +4)(y-3) ??

OpenStudy (anonymous):

how you do?

OpenStudy (anonymous):

I did by eye, u can also use quadratic formula.

OpenStudy (anonymous):

tell me quadratic formular please

OpenStudy (anonymous):

To do by eye, you want two numbers that add up to 1 multiply to -12.

OpenStudy (anonymous):

Quadratic formula: \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] for equation ax^2 + bx +c

OpenStudy (anonymous):

ok I saw it too

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Not finished yet...

OpenStudy (anonymous):

We have y = -4 or y = 3 and we have y = x^2 to get back to the original question.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So (x2 + 4)(x^2 -3) is your answer.

OpenStudy (anonymous):

ok I understand now

OpenStudy (anonymous):

OK.

OpenStudy (anonymous):

you do a good job

OpenStudy (anonymous):

np, anytime.